# The Representation Landscape of Few-Shot Learning and Fine-Tuning in Large Language Models

### v1

@@ 1 Introduction
    Both few-shot learning and fine-tuning show a clear division between model layers, a marked peak in the intrinsic dimension of the dataset, and a sudden change in the geometrical structure of the representations (see Sec. 4.1);
@@
    Few-shot learning leads to semantically meaningful clustering of the representations in early layers, organized hierarchically by subject (see Section 4.2);
@@
    Fine-tuning enhances the sharp emergence of clustered representations according to answers in the second half of the network (see Section 4.3).
@@ 3.2 Density peaks clustering
We run the experiments on a single Nvidia A100 GPU with 40GB memory. Extracting the hidden representation of 70 billion parameter models and their fine-tuning requires sharing the workload on 8 A100 GPUs. We provide code to reproduce our experiments at https://github.com/diegodoimo/geometry_icl_finetuning.
@@ 4 Results
We start by exploring the geometric properties of the representation landscape of LLMs. Our analysis proceeds from a broad description of the manifold geometry to its finer details. First, we measure the intrinsic dimension (ID) to understand the global structure of the data manifold. Next, we will describe the intermediate-scale behavior, counting the number of probability modes. Finally, we analyze the density distribution at the level of individual data points within clusters. These three quantities consistently show a two-phased behavior across the hidden layers of the LLMs we analyzed. All profiles of this and the following sections are smoothed using a moving average over two consecutive layers. We report the original profiles in the Appendix from Sec. D.2 to D.4 and in Sec. D.6.
@@ 5 Discussion and conclusion
We thank Alessandro Laio for many helpful discussions and valuable suggestions. We acknowledge the AREA Science Park supercomputing platform ORFEO and thank the technical support of the Laboratory of Data Engineering staff.
A.A., A.C., and D. D. were supported by the project “Supporto alla diagnosi di malattie rare tramite l’intelligenza artificial"- CUP: F53C22001770002. A.A., A. C. were supported by the European Union – NextGenerationEU within the project PNRR "PRP@CERIC" IR0000028 - Mission 4 Component 2 Investment 3.1 Action 3.1.1. A.S. was supported by the project PON “BIO Open Lab (BOL) - Raforzamento del capitale umano”- CUP: J72F20000940007.

### v2

@@ 1 Introduction
    Both few-shot learning and fine-tuning show a clear division between model layers, a marked peak in the intrinsic dimension of the dataset, and a sudden change in the geometrical structure of the representations (Secion 4.1);
@@
    Few-shot learning leads to semantically meaningful clustering of the representations in early layers, organized hierarchically by subject (Section 4.2);
@@
    Fine-tuning enhances the sharp emergence of clustered representations according to answers in the second half of the network (Section 4.3).
@@ 3.2 Density peaks clustering
We run the experiments on a single Nvidia A100 GPU with 40GB memory. Extracting the hidden representation of 70 billion parameter models requires 5 A100, and their fine-tuning requires 8 A100. We provide code to reproduce our experiments at https://github.com/diegodoimo/geometry_icl_finetuning.
@@ 4 Results
We start by exploring the geometric properties of the representation landscape of LLMs. Our analysis proceeds from a broad description of the manifold geometry to its finer details. First, we measure the intrinsic dimension (ID) to understand the global structure of the data manifold. Next, we will describe the intermediate-scale behavior, counting the number of probability modes on it. Finally, we analyze the density distribution at the level of individual data points within the clusters. These three quantities consistently show a two-phased behavior across the hidden layers of the LLMs we analyzed. All profiles of this and the following sections are smoothed using a moving average over two consecutive layers. We report the original profiles in the Appendix from Sec. D.2 to D.4 and in Sec. D.6.
@@ 5 Discussion and conclusion
We thank Alessandro Laio for many helpful discussions and valuable suggestions. We also thank the technical support of the Laboratory of Data Engineering staff and acknowledge the AREA Science Park supercomputing platform ORFEO.
A.A., A.C., and D. D. were supported by the project “Supporto alla diagnosi di malattie rare tramite l’intelligenza artificiale"- CUP: F53C22001770002. A.A., A. C. were supported by the European Union – NextGenerationEU within the project PNRR "PRP@CERIC" IR0000028 - Mission 4 Component 2 Investment 3.1 Action 3.1.1. A.S. was supported by the project PON “BIO Open Lab (BOL) - Raforzamento del capitale umano”- CUP: J72F20000940007.